# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v4

@@ Abstract
Current methodologies for novel view synthesis (NVS) in dynamic scenes encounter significant challenges in harmonizing memory consumption, model complexity, training efficiency, and rendering fidelity. Existing offline techniques, while delivering high-quality results, are often characterized by substantial memory demands and limited scalability. In contrast, online methods grapple with the challenge of balancing rapid convergence with model compactness. To address these issues, we propose continual dynamic neural graphics primitives (CD-NGP). Our approach synergizes features from both temporal and spatial hash encodings to achieve high rendering quality, employs parameter reuse to enhance scalability, and leverages a continual learning framework to mitigate memory overhead. Furthermore, we introduce a novel dataset comprising multi-view, exceptionally long video sequences with substantial rigid and non-rigid motion, thereby substantiating the scalability of our method.
We evaluate the quality, speed and scalability our method on both established public datasets and our exceptionally long video dataset. Notably, our method achieves an $85\%$ reduction in training memory consumption (less than 14GB) compared to offline techniques and significantly lowers streaming bandwidth requirements (less than 0.4MB/frame) relative to other online alternatives. The experimental results on our long video sequences dataset show the superior scalability and quality compared to existing state-of-the-art approaches.
@@
The demonstration of our continual learning pipeline for dynamic scenes. Our method segments input videos into multiple chunks, train the model chunk by chunk and leverage the re-use of parameters from the first branch to ensure high scalability. The spatial hash features could be encoded along voxels, orthogonally projected planes or both.
@@ Introduction
Neural radiance field (NeRF) [30] has made significant progress in 3D scene reconstruction and novel view synthesis. It takes multi-view RGB images with their camera poses as input and leverages multi-layer perceptrons (MLPs) to reconstruct both the geometry and color of the scene. Many variants have been proposed to improve NeRF’s efficiency [9, 38, 31, 56, 53], reconstruction quality [2, 3, 4], scalability [47, 50], and pose robustness [5, 26, 13, 52, 32, 51]. Recently, 3D Gaussian Splatting (3DGS) [22] proposes an explicit representation with differentiable rendering, which significantly speeds up the reconstruction and enables real-time novel view synthesis. Despite significant progress having been made in this field, the above-mentioned methods assume that the scene is static and may not work properly in dynamic scenes.
@@
To address the dynamic reconstruction problem, a number of efforts have been made in the offline setting, requiring access to the entire dataset during the training stage. For example, the methods in [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results. There are also methods speed up the optimization by using tensor factorization ([40, 17, 7]) and voxel representation [16, 45]. The more recent continual-learning-based methods [27, 48] can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference. However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the continual learning-based methods are weak in balancing the convergence speed and model compactness. As a result, these methods are limited in scalability, making them unsuitable for long videos.
In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos.
The proposed methodology is illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes: we partition dynamic videos into multiple segments based on timestamps and perform reconstruction using distinct model branches. Each branch includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$. For the first segment, we encode spatial features using a large hash table, and for subsequent segments, we use a smaller hash table. By reusing parameters from the initial branch, we enhance scalability and accelerate convergence. For each segment, we identify the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features using spatial hash grids. The spatial features from the initial branch and the current branch are fused through element-wise summation by default or concatenation if applicable. Temporal features are encoded by the hash table for the timestamp $t$ within the current branch only. These spatial and temporal features, along with the viewing direction $d$, are concatenated and input into MLP regressors for rendering and optimization.
We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and the dataset proposed by us. As illustrated in Fig. 1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly, in which we use metrics on the flame salmon scene in the DyNeRF dataset for fair comparison. The perceptual consistency is measured by $1-$ LPIPS by AlexNet [57, 24], and the speed is the reciprocal of the training time. The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
Figure 1: Comparison between different performance indexes of our method and recent representative baselines.
@@
Compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results. Compared with online baselines, our method achieves compact model size, high quality, and fast convergence simultaneously. Contributions of this paper are three-fold as follows:
@@
  • We propose a novel method combining spatial and temporal hash encodings for dynamic scenes and yield comparable results to both offline and online methods.
  • We propose a fast continual learning framework for dynamic scenes and reduce memory and storage occupation significantly.
@@
  • We provide a dataset of multi-view long video sequences ($>3$ minutes) with large motion and further demonstrate the scalability of our method on the proposed dataset.
@@ 1.1 Fast novel view synthesis in static scenes
The acceleration of NeRFs falls into two categories: training acceleration and rendering acceleration. Plenoxels [18] uses voxel grids to represent the scene and achieves magnitudes of acceleration over MLP backbones. TensoRF [9] decomposes the 3D scene as the product of tensors and achieves fast novel view synthesis. Plane-based methods [17, 21] project 3D points onto 3 orthogonal 2D planes to ensure sparsity and achieve fast convergence. Instant-NGP [31] leverages an occupancy grid and a multi-resolution hash encoding along with optimized tiny MLPs to enable fast novel view synthesis within a minute. The recent 3D Gaussian method [22] represents the scenes as a union of explicit Gaussian point clouds and achieves high fidelity, fast convergence, and real-time rendering on high-end hardware by eliminating the querying burden of spatial representations. KiloNeRF [37] distills the large MLP in NeRF [30] into tiny MLPs to accelerate rendering by thousands of times. Hedman et al. [20] propose the baking method (converting the MLP representation into explicit meshes) to accelerate the rendering process. They alpha-compose the cached features along the ray and use MLPs only once per ray to overcome the computation load of the per-point MLP querying in NeRF. BakedSDF [56] stores the view-dependent features into Gaussian spheres along the reconstructed surfaces, and achieves real-time rendering on ordinary laptops. MobileNeRF[10] represents scenes into meshes and enables real-time rendering on mobile devices. Light field methods [42, 8, 19] directly represent the scene as a set of rays to eliminate volume rendering and enable real-time applications on mobile devices.
@@ 1.2 Representations of dynamic scenes
There are mainly 3 variants of NeRF for dynamic scenes. The first class is pure MLP-based representation. D-NeRF [36] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes. DyNeRF [28] uses long latent vectors to encode temporal information along with large MLPs to achieve high-fidelity novel view synthesis in real dynamic scenes. Park et al. [34] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF.
The second class is tensor factorization-based representation. HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels. K-Planes [17] and HexPlane [7] factorize the whole 4D field into the product of 2D planes. Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on the DyNeRF dataset. The third class is the explicit voxel-based representations. TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation. It achieves high fidelity and fast convergence on the D-NeRF dataset. MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels. Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on DyNeRF dataset. Song et al. [43] train Instant-NGP [31] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes. Concurrent works are leveraging 4D Gaussian point clouds [55, 49] or 3D Gaussians with deformation networks [54] achieving real-time rendering and high fidelity in dynamic scenes.
@@ 1.3 Continual learning of NeRF
Continual learning (online learning) aims to adapt to new tasks while defying catastrophic forgetting (resisting performance downgrade on the previous tasks) [12]. There are mainly 3 approaches to achieving continual learning: replay, regularization, and parameter isolation. The replay-based methods either store the previous data in a buffer and replay them [39], or use the model to generate the pseudo-ground truth of previous tasks while training on new tasks [41]. The regularization-based methods hold the parameters close to the solution of previous tasks [14]. The parameter-isolation methods use different parameters to cope with different tasks. Recent efforts based on generative replay [6, 35, 11] enable continual learning of NeRF in static scenes and achieve high fidelity. StreamRF [27] applies continual NeRF to dynamic scenes by parameter isolation. Their explicit voxel representation enables fast reconstruction (75 minutes for 300 frames) and could be streamed frame by frame. INV [48] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation and achieves high-fidelity results in dynamic scenes.
@@ 2 Method
2.1 Preliminaries
Inspired by the previous NeRF variants [36, 28, 17, 7, 45, 31, 38] and continual learning methods [6, 35, 11, 27, 43, 48], we focus on the dynamic novel view synthesis problem in the multi-view setting and solve the problem in a continual learning manner to alleviate the huge memory consumption and bandwidth costs.
Following previous methods [36, 28, 17, 7], we consider the dynamic scene as a set of 3D points $\mathbf{p}=(x,y,z)$ with optical attributes $(\sigma,\mathbf{c})$, where $\sigma$ is the volume density of the point and $\mathbf{c}$ is the color of the point. We train our model along camera rays to reconstruct the scene. A camera ray is defined as:
@@ 2.1 Preliminaries
where $\mathbf{o}$ is the optical center of the camera, $\mathbf{d}$ is the direction off the camera. For each point along the camera ray with coordinates $(x,y,z)$ observed at direction $\mathbf{d}=(\phi,\psi)$, a field function $F$ is queried to obtain the optical attributes $(\sigma,\mathbf{c})$. Conditioned on time $t$, the entire model is defined as:
@@
[…]
To simplify the notations, let the whole sequence of frames be evenly divided into $k$ chunks where a chunk $\mathbf{T_{k}}=\{\tau_{i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \lfloor\frac{N_{time}}{T_{chunk}}\rfloor=k\}$ is a sequential subset of length $T_{chunk}$. We define the continual learning problem as learning a representation where only the frames in the current chunk are available at each optimization step, i.e. $\mathbf{T_{a}}=\mathbf{T_{k}}$. Accordingly, $T_{chunk}=1$ denotes training the model online in a frame-by-frame manner, $T_{chunk}=10$ denotes training the model online with a buffer of 10 frames. If the video contains 300 frames in total, the $T_{chunk}=300$ denotes training the entire 300-frame sequence offline. The definition is close to real-world applications since caching all training frames in memory or storing them on the disk is limited in scalability for long videos.
@@ 2.2 Representation of short dynamic intervals
As shown in [46], naive 4D hash grids for dynamic scenes are short in reconstruction quality. To achieve fast reconstruction and slim model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion. The spatial hash features may be encoded along 3D voxels [31], 2D planes [17, 21], or both [38]. The time feature is encoded by short hash tables in a per-frame manner. Following previous state-of-the-art methods [30, 31, 16], we apply a two-fold field function for geometry and color individually. The architecture of the field function is thus defined as:
@@
$\mathbf{c}=f_{\theta_{2}}(\mathbf{H},\phi,\psi).$ (4)
where $\mathbf{x}=(x,y,z)$ is the spatial coordinates, $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, and $\oplus$ is the concatenation of vectors. The MLP $f_{{\theta_{1}}}$ in Eq. 3 fuses features across the spatial domain and time domain and obtains the volume density $\sigma$. View-dependent colors $\mathbf{c}$ are regressed by another tiny MLP with parameter $\theta_{2}$ in Eq. 4. The spatial encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)$ (5)
where the $P_{m}(x,y,z)\in\mathbf{R}^{F}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t. 3D coordinates $(x,y,z)$ at the $l$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31]. If necessary, Eq. 5 could be adopted to the hybrid (MERF) representation [38], defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (6)
where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t. axes $(i,j)$ at the $m$-th resolution. Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].

### v5

@@ Abstract
Current methods for novel view synthesis (NVS) in dynamic scenes encounter significant challenges in managing memory consumption, model complexity, training efficiency, and rendering fidelity. Existing offline techniques, while delivering high-quality results, face challenges from substantial memory demands and limited scalability. Conversely, online methods struggle to balance rapid convergence with model compactness. To address these issues, we propose continual dynamic neural graphics primitives (CD-NGP). Our approach leverages a continual learning framework to reduce memory overhead, and it also integrates features from distinct temporal and spatial hash encodings for high rendering quality. Meanwhile, our method employs parameter reuse to achieve high scalability. Additionally, we introduce a novel dataset featuring multi-view, exceptionally long video sequences with substantial rigid and non-rigid motion, which is seldom possessed by existing datasets. We evaluate the reconstruction quality, speed and scalability of our method on both the established public datasets and our exceptionally long video dataset. Notably, our method achieves an $85\%$ reduction in training memory consumption (less than 14GB) compared to offline techniques and significantly lowers streaming bandwidth requirements (less than 0.4MB/frame) relative to other online alternatives. The experimental results on our long video sequences dataset show the superior scalability and reconstruction quality compared to existing state-of-the-art approaches.
@@
The outline of our continual learning pipeline for dynamic scenes. Our method segments input videos into multiple chunks and encodes the chunks by separate spatial and time hash encodings. Subsequently, we train the auxiliary branches incrementally on each chunk and leverage the re-use of parameters in the base branch to ensure high scalability. The spatial hash features could be encoded according to voxel grids, orthogonally projected plane grids or a combination of both.
@@ Introduction
Neural radiance field (NeRF) introduced in [30] has made significant progress in 3D scene reconstruction and novel view synthesis. It takes multi-view RGB images with their camera poses as input and leverages multi-layer perceptrons (MLPs) to reconstruct both the geometry and color of the scene. Many variants have been proposed to improve NeRF’s efficiency [9, 38, 31, 56, 53], reconstruction quality [2, 3, 4], scalability [47, 50], and pose robustness [5, 26, 13, 52, 32, 51]. Recently, 3D Gaussian Splatting (3DGS) [22] proposes an explicit representation with differentiable rendering, which significantly speeds up the reconstruction and enables real-time novel view synthesis. Despite significant progress having been made in this field, the above-mentioned methods assume that the scene is static and may not work properly in dynamic scenes.
@@
To address the dynamic reconstruction problem, a number of efforts have been made in the offline setting, requiring access to the entire dataset during the training stage. For example, the methods in [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results. There are also methods that speed up the optimization by using the tensor factorization ([40, 17, 7]) and voxel representations [16, 45]. The more recent continual-learning-based methods [27, 48] (also mentioned as online methods) can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference. However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the online methods are weak in balancing the convergence speed and model compactness. These limitations in scalability makes them impractical for the wide application on long videos.
In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos. As illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes, CD-NGP derives from three key ideas: dividing the video into segments, encoding spatial and temporal features separately, and reusing features across different segments to ensure high scalability: Firstly, we partition dynamic videos into multiple segments according to time stamps $t$ and perform reconstruction using distinct model branches to facilitate continual learning. Each branch of our model includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$. Secondly, the spatial and temporal features are encoded separately to avoid feature confusion in different axes. For each model branch, we sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids like [31], and encode the points into spatial features using spatial hash grids. We also encode temporal features using hash grids for the separated $t$ axis. These spatial and temporal features and the viewing direction $\mathbf{d}$ are concatenated and input into MLP regressors for rendering and optimization. Thirdly, we leverage the redundancy in the dynamic scenes and fuse the spatial features of different branches to yield high scalability. We set a larger hash table size ($2^{19}$) in the initial branch (base branch), and set a smaller size ($2^{14}$) for the other branches (auxiliary branches). This asymetric design facilitates reusing the features trained in the base branch, thereby reducing storage consumption and boosting rendering quality.
Figure 1: Comparison between different performance indexes of our method and recent representative baselines. The perceptual consistency is measured by $1-$ LPIPS by AlexNet [57, 24], and the speed is the reciprocal of the training time. The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
@@
We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and a self-collected dataset. As illustrated in Fig. 1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly. Compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results. Compared with online baselines, our method achieves compact model size, high rendering quality, and fast convergence simultaneously. The contributions of this paper are three-fold as follows:
@@
  • We propose a novel method combining separate spatial and temporal hash encodings, achieving high scalability and rendering quality in dynamic scenes.
  • We propose a fast and highly scalable continual learning framework for dynamic scenes. It enables reconstructing long videos with low memory usage and low storage consumption. Our method also outperforms online alternatives and achieves comparable results to offline methods.
@@
  • We provide a dataset of multi-view exceptionally long video sequences ($>3$ minutes) with large non-rigid motion, which is seldom possessed by existing datasets.
@@ 1.1 Fast novel view synthesis in static scenes
The acceleration of NeRFs falls into two categories: training acceleration and rendering acceleration. Plenoxels [18] uses voxel grids to represent the scene and achieves magnitudes of acceleration over MLP backbones. TensoRF [9] decomposes the 3D scene as the product of tensors and achieves fast novel view synthesis. Plane-based methods [17, 21] project 3D points onto 3 orthogonal 2D planes to ensure sparsity and achieve fast convergence. Instant-NGP [31] leverages an occupancy grid and a multi-resolution hash encoding along with optimized tiny MLPs to enable fast novel view synthesis within a minute. The recent 3D Gaussian method [22] represents the scenes as a union of explicit Gaussian point clouds and achieves high fidelity, fast convergence, and real-time rendering on high-end hardware by eliminating the querying burden of spatial representations. KiloNeRF [37] distills the large MLP in NeRF [30] into tiny MLPs to accelerate rendering by thousands of times. Hedman et al. [20] propose the baking method (converting the MLP representation into explicit meshes) to accelerate the rendering process. They alpha-compose the cached features along the ray and use MLPs only once per ray to overcome the computation load of the per-point MLP querying in NeRF. BakedSDF [56] stores the view-dependent features into Gaussian spheres along the reconstructed surfaces, and it achieves real-time rendering on ordinary laptops. MobileNeRF[10] represents scenes into meshes and enables real-time rendering on mobile devices. Light field methods [42, 8, 19] directly represent the scene as a set of rays to eliminate volume rendering and enable real-time applications on mobile devices.
@@ 1.2 Representations of dynamic scenes
There are mainly 3 variants of NeRF for dynamic scenes. The first class is pure MLP-based representation. D-NeRF [36] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes. DyNeRF [28] uses long latent vectors to encode temporal information along with large MLPs to achieve high-fidelity novel view synthesis in real dynamic scenes. Park et al. [34] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF. The second class is tensor factorization-based representation. HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels. K-Planes [17] and HexPlane [7] factorize the whole 4D field into the product of 2D planes. Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on the DyNeRF dataset. The third class is the explicit voxel-based representations. TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation, and it achieves high fidelity and fast convergence on the D-NeRF dataset. MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels. Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on DyNeRF dataset. Song et al. [43] train Instant-NGP [31] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes. Concurrent works are leveraging 4D Gaussian point clouds [55, 49] or 3D Gaussians with deformation networks [54] achieving real-time rendering and high fidelity in dynamic scenes.
@@ 1.3 Continual learning of NeRF
Continual learning (online learning) aims to adapt to new tasks while defying catastrophic forgetting (resisting performance downgrade on the previous tasks) [12]. There are mainly 3 approaches to achieving continual learning: replay, regularization, and parameter isolation. The replay-based methods either store the previous data in a buffer and replay them [39], or use the model to generate the pseudo-ground truth of previous tasks while training on new tasks [41]. The regularization-based methods hold the parameters close to the solution of previous tasks [14]. The parameter-isolation methods use different parameters to cope with different tasks. Recent efforts based on generative replay [6, 35, 11] enable continual learning of NeRF in static scenes and achieve high fidelity. StreamRF [27] applies continual NeRF to dynamic scenes by parameter isolation. Their explicit voxel representation enables fast reconstruction (75 minutes for 300 frames) and could be streamed frame by frame. INV [48] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation. It achieves high reconstruction quality in dynamic scenes.
@@ 2 Method
2.1 Problem definition and preliminaries
[…]
Inspired by the previous NeRF variants [36, 28, 17, 7, 45, 31, 38] for dynamic novel view synthesis and continual learning methods [6, 35, 11, 27, 43, 48] in the static scenes, we focus on the dynamic novel view synthesis problem in a multi-view setting. We solve the problem using continual learning to alleviate the huge memory consumption and bandwidth costs. Unlike the existing replay-based continual learning methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation. This is because the current offline state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are regressions of the optical attributes in 3D space within a limited time instead of the object trajectory and inherent dynamics. Therefore, the replay method is of limited use in our defined scenario: Since new transient processes are constantly arriving, the model of a limited size is not capable of reconstructing unlimited transient effects. Thus, the replay-based method suffers from catastrophic forgetting. The design of our model follow the parameter isolation approach: Let the time stamps be divided into $N_{chunk}$ chunks, the model consists of $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly.
Following previous methods [36, 28, 17, 7], we consider the dynamic scene as a set of 3D points $\mathbf{x}=(x,y,z)$ with optical attributes $(\sigma,\mathbf{c})$, where $\sigma$ is the volume density of the point and $\mathbf{c}$ is the color of the point. To reconstruct the scene, we train our model along camera rays which are defined as:
@@ 2.1 Problem definition and preliminaries
where $\mathbf{o}$ is the optical center of the camera, $\mathbf{d}$ is the direction from the camera to the scenes. For each point along the camera ray with coordinates $\mathbf{x}=(x,y,z)$ observed at direction $\mathbf{d}=(\phi,\psi)$, a field function $F$ is queried to obtain the optical attributes $(\sigma,\mathbf{c})$. Conditioned on time $t$, the entire model is defined as:
@@ 2.2 Representation of short dynamic intervals
To reconstruct the individual chunks of the dynamic scenes, we propose a representation of short dynamic intervals, which serves as the structures of the different branches of our model and handle the dynamic scnes in one chunk efficiently. As shown in [46], naive 4D hash grids for dynamic scenes are short in reconstruction quality. To achieve fast reconstruction and compact model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion. The spatial hash features can be encoded along 3D voxels [31], 2D planes [17, 21], or both [38]. The time feature is encoded by short hash tables in a per-frame manner. Following previous state-of-the-art methods [30, 31, 16], we apply a two-fold field function for geometry and color individually. The architecture of the field function is thus defined as:
@@
$\mathbf{c}=f_{\theta_{2}}(\mathbf{H},\phi,\psi),$ (4)
where $\mathbf{x}=(x,y,z)$ is the spatial coordinates, $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, and $\oplus$ is the concatenation of vectors. The MLP $f_{{\theta_{1}}}$ in Eq. 3 fuses features across the spatial domain and time domain and obtains the volume density $\sigma$. View-dependent colors $\mathbf{c}$ are regressed by another tiny MLP with parameter $\theta_{2}$ in Eq. 4. The spatial hash encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z),$ (5)
where the $P_{m}(x,y,z)\in\mathbf{R}^{F}$ denotes the tri-linearly interpolated feature from the nearby 8 grid vertices w.r.t. 3D coordinates $(x,y,z)$ at the $m$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31]. The voxel-based hash encoding structure is flexible and could be replaced by other spatial hash encodings, which will be discussed in the experiments.